Find the equation of the line through the point (1,2) whose distance from the point(3, 1) is √5

Grade:11

1 Answers

Waseem khan
6 years ago
From your question we see that the point (1,2) and (3, 1) are at the distance of √5 so the slope of the line passing through this point is - 1/2, so the slope of the line passing through (1,2) will be 2, so the equation will be (Y-2)=2(X-1), SO FROM HERE WE GET 2X-Y=0